OpenStudy (anonymous):

g(x)=x^2+2x-6 find the value of 2[g(2)] 6 4 18 12

OpenStudy (anonymous):

plug in 2 for x and then multiply by 2 what do you think the answer is?

OpenStudy (anonymous):

is six the answer? or did i do it wrong

OpenStudy (anonymous):

(2)^2 + 2(2) - 6 4 + 4 - 6 = ?

OpenStudy (anonymous):

4+4=8-6=2

OpenStudy (anonymous):

perfect - now multiply that by 2

OpenStudy (anonymous):

2x2=4...

OpenStudy (anonymous):

you got it :)

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

You are welcome